# Comparing cosmic shear nulling methods for Stage-IV surveys

### v1

@@ 1 Introduction
Cosmic shear measures the clustering of the large-scale-structure (LSS) from the highly non-linear and non-Gaussian sub-Mpc regime, out to much larger linear scales. By measuring the correlation of galaxy shapes at and between different redshifts, the evolution of the LSS can be traced, enabling us to detect the effect of dark energy on the growth of structure. Since gravitational lensing is not sensitive to the dynamical state of the intervening masses, it yields a direct measure of the total mass, which is predominantly dark matter. As such, cosmic shear is now a well established cosmological tool, with stage III surveys providing tight constraints on the matter density parameter $\Omega_{\rm m}$ and the amplitude of matter fluctuations $\sigma_{8}$ (amon/etal:2022; dalal/etal:2023; wright/etal:2025). More specifically cosmic shear is most sensitive to the combination of these two parameters given by $S_{8}=\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$. Photometric galaxy surveys, such as those performed by the Euclid space telescope (euclid/mellier:2024) and Rubin observatory (mandelbauam/etal:2018), will soon be observing almost the entire observable extragalactic sky, with cosmic shear measurements being a central part of their scientific program. These surveys will therefore benefit from having access to higher statistical power at the larger scales which are easier to model. Conversely this means that to gain any further constraining power from weak lensing data sets will require including smaller scales where modelling the non-linear signal is a known challenge (for example, preston/etal:2023), in particular baryon feedback poses a large uncertainty at these scales. While the constraining power for current surveys has meant that current methods have been sufficient, predictions for upcoming analyses from Euclid and Rubin observatory have forecast severe scale cuts (an $\ell_{\rm max}=1500$ or $\ell_{\rm max}=2000$ depending on the forecast) would be required to recover unbiased cosmological constraints (martinelli/etal:2021; huang/etal:2019), whilst alternative approaches to modelling these scales with the necessarily larger parameter space leads to a huge decrease in precision (semboloni/etal:2013; huang/etal:2019; spuriomancini/bose:2023; boruah/etal:2024).
@@
Baryon feedback generally refers to any process by which the dynamics of baryons present influences that of the wider matter distribution. This includes feedback processes from Active Galactic Nuclei which cause gas to be ejected from massive haloes, and supernovae and star formation processes, which are the dominant effects at $k\gtrsim 10{\rm Mpc}^{-1}$. These processes suppress clustering at small scales and therefore measurements of cosmic shear are impacted. Cosmological hydrodynamical simulations have been able to establish predictions for how baryon feedback processes suppress the matter power spectrum (schaye/etal:2010; vandaalen/etal:2020; schaye/etal:2023; salcido/etal:2023; pakmor/etal:2023; mccarthy/etal:2023; schaller/etal:2025), however the predicted amplitude of this suppression varies from less than 10% to more than 30% around $k=4\,{\rm Mpc}^{-1}$, and can still be up to 10% at $k=0.3\,{\rm Mpc}^{-1}$, depending on the simulation. Furthermore, simulations with large enough volumes relevant for clustering studies are unable to resolve the small scales required to fully include feedback processes. To get around this, these processes are included via the sub-grid physics, however it has been shown that particular simulation observables are susceptible to the details of these prescriptions. Incorrect specification of these effects will bias the resulting matter power spectrum prediction, and therefore the inferred cosmological parameters as well (e.g., 2018MNRAS.480.2247C; 2020MNRAS.493.1640C).
@@
Previous analyses have linked the suppression of small scale matter clustering to the baryon fraction of groups and clusters (semboloni/etal:2013; vandaalen/etal:2020; salcido/etal:2023; salcido/etal:2025). As such, observations that are sensitive to the gas distribution have been used to constrain the impact of baryon feedback or calibrate simulations. X-ray observations, which are sensitive to the hot gas inside galaxies and clusters, have been combined with other probes that measure the total mass, like weak lensing, to estimate the hot gas fraction (for example sun/etal:2009; lovisari/etal:2015; eckert/etal:2016; akino/etal:2022; ferreira/etal:2024; laposta/etal:2024). Alternatively, measurements of the Sunyaev-Zel’dovich (SZ) effect from CMB experiments have been shown to be able to constrain baryon feedback, either alone (bolliet/etal:2018; reichardt/etal:2021; efstathiou/etal:2025) or in combination with weak lensing (amodeo/etal:2021; troester/etal:2022; pandey/etal:2023; mccarthy/etal:2023; to/etal:2024; bigwood/etal:2024; mccarthy/etal:2024; hadzhiyska/etal:2024; mccarthy/etal:2024). Data from probes that trace baryonic effects directly will potentially be sufficient for surveys to achieve the accuracy and percent-level precision for which they have been designed, however wayland/etal:2025 predict that for this to be achievable a CMB-S4-like survey or X-ray data from around 5000 clusters would be required, which are both several years away.
@@
The contamination of weak lensing measurements from small-scale uncertainties has motivated analysis teams to remove correlations at small angular scales from their data vectors (2021arXiv210513548K; 2025arXiv251005539E). This approach has the disadvantage that it does not cleanly remove the small physical (3D) scales where baryon feedback is significant, due to the fact that a given angular scale receives contributions from a wide range of physical scales, a consequence of the broad radial lensing kernel. Weak lensing measures the 2D projection of the matter distribution, meaning large-scale fluctuations from distant lenses contribute to the same angular scales as small-scale fluctuations from nearby lenses. Applying cuts in angular space, e.g., via a maximum harmonic-space multipole $\ell$, is therefore not optimal.
@@
In this work, we conduct an assessment of the performance of these three approaches in the context of the forthcoming Stage-IV cosmic shear surveys that aim to constrain Dark Energy physics at the percent level. As part of this exercise, we develop the ‘$k$-cut’ method presented in hutererandwhite:2005, which to our knowledge has not been considered in the literature since it was first proposed. We also adapt the cross-correlation method from the CMB lensing literature and apply it to cosmic shear in combination with photometric galaxy clustering.
@@ 2 Cosmic Shear Theory
In this section we briefly introduce some cosmic shear formalism (see kilbinger/etal:2017 for further details). For a pair of tomographic bins, $i$ and $j$, an accurate expression for the cosmic shear power spectrum, valid in the Limber approximation, is given by
@@ 3.1 LU nulling
We describe here a new implementation of a method first proposed in hutererandwhite:2005, which to our knowledge has not been since considered in the literature. We will hereafter refer to this method as ‘LUnul’, and provide a public implementation [^1] https://github.com/ahallcosmo/lunul.
@@
where $\mathbf{c}_{\ell}$ is the vector of angular power spectra for each tomographic bin pair at a given $\ell$; if there are $N$ tomographic bins, this vector has length $N_{p}=N(N+1)/2$. The matrix $\mathbf{P}_{\ell}$ has dimensions $N_{p}\times N_{k}$, and $\mathbf{1}$ is a vector of ones with length $N_{k}$. This matrix formulation of the Limber integral has some computational advantages, with the drawback that $N_{k}$ is required to be the same for every bin pair.
@@
Our strategy is therefore to define a $k_{\rm max}$ and, for each $\ell$, discard rows from $\mathbf{L}_{\ell}^{-1}\mathbf{c}_{\ell}$ that have contribution from $k>k_{\rm max}$. If no such rows are available, we discard all of the rows, which is equivalent to throwing out the whole data vector for that $\ell$. Since higher $k$ contribute progressively more to higher $\ell$ in a smooth fashion, this procedure results in a softer form of the standard scale cut method.
There is some freedom in the implementation of the LUnul method. Firstly, there is sensitivity to the way in which the Limber integral is discretized, especially because only the first $N_{p}$ modes can formally be discarded and typically $N_{p}\ll N_{k}$. We found a linear spacing in $\log k$ works well, having also experimented with Monte Carlo sampling of the Limber integrand and the list of nodes points returned by the scipy.integrate.quad routine for a fixed choice of bin pair. These latter two choices of nodes can work well for bin pairs close to the fiducial bin pair that generated them, but become suboptimal for other bin pairs (the disadvantage of enforcing a constant $k$-sampling for every bin pair, as referenced earlier). We also introduce a tolerance parameter, which allows rows of $\mathbf{U}_{\ell}$ to be retained if modes having $k>k_{\rm max}$ contribute less than a fraction tol to the total over all $k$. The results in this paper set tol=0.01, which gives excellent results. We also have the freedom to choose the number of nodes, $N_{k}$, used to discretize the integral. In general, this should be large enough to attain accuracy in the power spectrum within some chosen tolerance, but small enough that the method retains some of its unique features; if $N_{k}$ is too large, then the first $N_{p}$ rows of the transformed data vector will only have the highest $k$ nodes nulled out, and the method will essentially be equivalent to discarding $\ell$-modes that have more than a fraction tol contribution from $k>k_{\rm max}$. We have found $N_{k}=\mathcal{O}(100)$ works well.
@@ 3.2 BNT
The BNT transform (bnt:2014) is a linear transform which reorganises the cosmic shear signal by separating the contribution from foreground structure at different redshifts. The transformed lensing kernels are narrower and as such the resulting power spectra has a closer mapping between $k$ and $\ell$. All information is preserved since the BNT transform is linear and invertible. This method benefits from a large number of tomographic bins and therefore has had limited impact of current Stage III surveys (taylor/etal:2021) but has been shown to be able to significantly improve constraining power for a Euclid-like survey (gu/etal:2025).
The BNT-transformed shear signal is given by
$\hat{\gamma}^{a}=\sum_{i=1}^{N}f_{i}^{a}\gamma^{i},$ (6)
with $\hat{q}^{a}(\chi)$ the corresponding transformed lensing window function. $N$ here is the number of tomographic bins. The transformation only acts on the tomographic bin space (mapping indices $i$ to transformed indices $a$), leaving the angular space unchanged. The transformed shear power spectrum is denoted $\hat{C}_{\ell,\gamma\gamma}^{ab}$.
@@
The advantage of the BNT transform is that it operates directly on the shear signal, and hence can be useful for any statistics based on the shear signal, not just the two-point statistics. It should be noted, however, that simply re-weighting the maps is not sufficient as it couples the noise, which is problematic for some high-order statistics. Methods have been developed to account for this, by applying the BNT method, smoothing the maps and then transforming back to the original basis as implemented in 2025arXiv251004953E. Another advantage is that the transformation is linear and invertible, and hence lossless.
After transformation, the mapping between angular space and 3D physical space is much cleaner in the power spectrum. Subsequent cuts in $\ell$ on the transformed power spectra can hence provide a cleaner separation of non-linear scales (taylor/etal:2018; 2020PhRvD.102h3535D; taylor/etal:2021). The performance of this method is most effective when many tomographic bins are available. An example of the impact of the BNT transformation on the lensing kernels is described in Section 4.
@@ 3.3 Cross-correlation
CMB studies have used cross-correlations with foreground galaxy samples to remove the lensing contribution from the matter distribution at the redshifts of these galaxies (e.g., qu/etal:2023). The aim in those analyses was to isolate lensing due to structure at a higher redshift, but its application to cosmic shear holds potential benefits by removing the low redshift contribution. At a fixed angular scale, it is the baryon suppression at low redshifts (equivalent to small physical scales) that contaminates the cosmic shear signal from higher redshift (equivalent to larger, more linear scales). Tracers at lower redshift can hence be used to partially remove the contribution to the shear signal from these low-redshift, small-scale structures.
@@ 4 Mock Data
The cross-correlation method relies on a low-redshift lens sample used to ‘clean’ a higher redshift source sample. When comparing the performance of this method against other methods that use only the source sample, we must be careful to account for the additional information brought by the lens density field. To perform as close an ‘apples-to-apples’ comparison as possible, we fix the total (i.e., lens + source) sample and assume that the lowest two source redshift bins can be repurposed as lens bins. The third source bin has strong overlap with neighbouring bins and is not expected to benefit significantly from the cross-correlation nulling, so we discard it. This configuration of two low-redshift lens bins and three high-redshift source bins should be roughly optimal for this form of nulling, but we caution against overinterpretation of the comparison with the other nulling methods. As the cross-correlation method has not been considered in the context of 3x2pt samples before, we consider this setup sufficient as a first test of its performance. The lensing kernels shown in Figure 2, for the standard case in the top panel and the BNT-transformed case in the lower panel.
@@
Baryon feedback is included using HMCode2020 (mead/etal:2021) via the parameter $\Theta_{\rm AGN}$ which modulates the strength of the baryon feedback suppression. This is not a physical parameter but is related to an AGN heating parameter in the BAHAMAS simulations, which HMCode2020 was designed to be able to recover. We use $\Theta_{\rm AGN}=7.8$ which matches the fiducial BAHAMAS simulation – the ratio of the matter power spectrum with and without baryon feedback is shown in Figure 3 for a range of redshift values.
Figure 3: Suppression of the matter power spectrum due to baryon feedback ($\Theta_{\rm AGN}=7.8$) at different redshifts.

### v2

@@ 1 Introduction
Cosmic shear measures the clustering of the large-scale-structure (LSS) from the highly non-linear and non-Gaussian sub-Mpc regime, out to much larger linear scales. By measuring the correlation of galaxy shapes at and between different redshifts, the evolution of the LSS can be traced, enabling us to detect the effect of dark energy on the growth of structure. Since gravitational lensing is not sensitive to the dynamical state of the intervening masses, it yields a direct measure of the total mass, which is predominantly dark matter. As such, cosmic shear is now a well established cosmological tool, with stage III surveys providing tight constraints on the matter density parameter $\Omega_{\rm m}$ and the amplitude of matter fluctuations $\sigma_{8}$ (Amon et al., 2022; Dalal et al., 2023; Wright et al., 2025). More specifically cosmic shear is most sensitive to the combination of these two parameters given by $S_{8}=\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$. Photometric galaxy surveys, such as those performed by the Euclid space telescope (Euclid Collaboration et al., 2024) and Rubin observatory (The LSST Dark Energy Science Collaboration: Mandelbaum et al., 2018), will soon be observing almost the entire observable extragalactic sky, with cosmic shear measurements being a central part of their scientific program. These surveys will therefore benefit from having access to higher statistical power at the larger scales which are easier to model. Conversely this means that to gain any further constraining power from weak lensing data sets will require including smaller scales where modelling the non-linear signal is a known challenge (for example, Preston et al., 2023), in particular baryon feedback poses a large uncertainty at these scales. While the constraining power for current surveys has meant that current methods have been sufficient, predictions for upcoming analyses from Euclid and Rubin observatory have forecast severe scale cuts (an $\ell_{\rm max}=1500$ or $\ell_{\rm max}=2000$ depending on the forecast) would be required to recover unbiased cosmological constraints (Martinelli et al., 2021; Huang et al., 2019), whilst alternative approaches to modelling these scales with the necessarily larger parameter space leads to a huge decrease in precision (Semboloni et al., 2013; Huang et al., 2019; Spurio Mancini and Bose, 2023; Boruah et al., 2024).
@@
Baryon feedback generally refers to any process by which the dynamics of baryons present influences that of the wider matter distribution. This includes feedback processes from Active Galactic Nuclei which cause gas to be ejected from massive haloes, and supernovae and star formation processes, which are the dominant effects at $k\gtrsim 10{\rm Mpc}^{-1}$. These processes suppress clustering at small scales and therefore measurements of cosmic shear are impacted. Cosmological hydrodynamical simulations have been able to establish predictions for how baryon feedback processes suppress the matter power spectrum (Schaye et al., 2010; van Daalen et al., 2020; Schaye et al., 2023; Salcido et al., 2023; Pakmor et al., 2023; McCarthy et al., 2023; Schaller and Schaye, 2025), however the predicted amplitude of this suppression varies from less than 10% to more than 30% around $k=4\,{\rm Mpc}^{-1}$, and can still be up to 10% at $k=0.3\,{\rm Mpc}^{-1}$, depending on the simulation. Furthermore, simulations with large enough volumes relevant for clustering studies are unable to resolve the small scales required to fully include feedback processes. To get around this, these processes are included via the sub-grid physics, however it has been shown that particular simulation observables are susceptible to the details of these prescriptions (e.g., Debackere et al. 2020; Castro et al. 2021). Incorrect specification of these effects will bias the resulting matter power spectrum prediction, and therefore the inferred cosmological parameters as well (e.g., Copeland et al. 2018, 2020).
@@
Previous analyses have linked the suppression of small scale matter clustering to the baryon fraction of groups and clusters (Semboloni et al., 2013; van Daalen et al., 2020; Salcido et al., 2023; Salcido and McCarthy, 2025). As such, observations that are sensitive to the gas distribution have been used to constrain the impact of baryon feedback or calibrate simulations. X-ray observations, which are sensitive to the hot gas inside galaxies and clusters, have been combined with other probes that measure the total mass, like weak lensing, to estimate the hot gas fraction (for example Sun et al., 2009; Lovisari et al., 2015; Eckert et al., 2016; Akino et al., 2022; Ferreira et al., 2024; La Posta et al., 2024). Alternatively, measurements of the Sunyaev-Zel’dovich (SZ) effect from CMB experiments have been shown to be able to constrain baryon feedback, either alone (Bolliet et al., 2018; Reichardt et al., 2021; Efstathiou and McCarthy, 2025) or in combination with weak lensing (Amodeo et al., 2021; Tröster et al., 2022; Pandey et al., 2023; McCarthy et al., 2023; To et al., 2024; Bigwood et al., 2024; McCarthy et al., 2024; Hadzhiyska et al., 2024; McCarthy et al., 2024). Data from probes that trace baryonic effects directly will potentially be sufficient for surveys to achieve the accuracy and percent-level precision for which they have been designed, however Wayland et al. (2025) predict that for this to be achievable a CMB-S4-like survey or X-ray data from around 5000 clusters would be required, which are both several years away.
@@
The contamination of weak lensing measurements from small-scale uncertainties has motivated analysis teams to remove correlations at small angular scales from their data vectors (Krause et al., 2021; Emas et al., 2025). This approach has the disadvantage that it does not cleanly remove the small physical (3D) scales where baryon feedback is significant, due to the fact that a given angular scale receives contributions from a wide range of physical scales, a consequence of the broad radial lensing kernel. Weak lensing measures the 2D projection of the matter distribution, meaning large-scale fluctuations from distant lenses contribute to the same angular scales as small-scale fluctuations from nearby lenses. Applying cuts in angular space, e.g., via a maximum harmonic-space multipole $\ell$, is therefore not optimal.
@@
In this work, we conduct an assessment of the performance of these three approaches in the context of the forthcoming Stage-IV cosmic shear surveys that aim to constrain Dark Energy physics at the percent level. As part of this exercise, we develop the ‘$k$-cut’ method presented in Huterer and White (2005), which to our knowledge has not been considered in the literature since it was first proposed. We also adapt the cross-correlation method from the CMB lensing literature and apply it to cosmic shear in combination with photometric galaxy clustering.
@@ 2 Cosmic Shear Theory
In this section we briefly introduce some cosmic shear formalism (see Kilbinger et al. 2017 for further details). For a pair of tomographic bins, $i$ and $j$, an accurate expression for the cosmic shear power spectrum, valid in the Limber approximation, is given by
@@ 3.1 LU nulling
We describe here a new implementation of a method first proposed in Huterer and White (2005), which to our knowledge has not been since considered in the literature. We will hereafter refer to this method as ‘LUnul’, and provide a public implementation [^1] https://github.com/ahallcosmo/lunul.
@@
where $\mathbf{c}_{\ell}$ is the vector of angular power spectra for each tomographic bin pair at a given $\ell$; if there are $N$ tomographic bins, this vector has length $N_{p}=N(N+1)/2$. The matrix $\mathbf{P}_{\ell}$ has dimensions $N_{p}\times N_{k}$, and $\mathbf{1}$ is a vector of ones with length $N_{k}$. This matrix formulation of the Limber integral has some computational advantages, with the potential to reduce computational time, however the drawback that $N_{k}$ is required to be the same for every bin pair.
@@
Our strategy is therefore to define a $k_{\rm max}$ and, for each $\ell$, discard rows from $\mathbf{L}_{\ell}^{-1}\mathbf{c}_{\ell}$ that have contribution from $k>k_{\rm max}$. If no such rows are available, we discard all of the rows, which is equivalent to throwing out the whole data vector for that $\ell$. Since higher $k$ contribute progressively more to higher $\ell$ in a smooth fashion, this procedure results in a softer form of the standard scale cut method. Effectively this method is not removing individual tomographic spectra, but instead forming linear combinations of them that cancel specific parts of the Limber integrand. Since high-$k$ power mostly comes from low-redshift structure, where all lensing kernels look similar, the LU decomposition results in the common high-$k$ part being cancelled out, while the differences at lower-$k$ (which are dependent on the redshift/bin pair) are retained.
There is some freedom in the implementation of the LUnul method. Firstly, there is sensitivity to the way in which the Limber integral is discretized, especially because only the first $N_{p}$ modes can formally be discarded and typically $N_{p}\ll N_{k}$. We found a linear spacing in $\log k$ works well, having also experimented with Monte Carlo sampling of the Limber integrand and the list of nodes points returned by the scipy.integrate.quad routine for a fixed choice of bin pair. These latter two choices of nodes can work well for bin pairs close to the fiducial bin pair that generated them, but become suboptimal for other bin pairs (the disadvantage of enforcing a constant $k$-sampling for every bin pair, as referenced earlier). We also introduce a tolerance parameter, which allows rows of $\mathbf{U}_{\ell}$ to be retained if modes having $k>k_{\rm max}$ contribute less than a fraction tol to the total over all $k$. The results in this paper set tol=0.01, which gives excellent results. We considered a range a different values for the tolerance parameter and found the same trends. We also have the freedom to choose the number of nodes, $N_{k}$, used to discretize the integral. In general, this should be large enough to attain accuracy in the power spectrum within some chosen tolerance, but small enough that the method retains some of its unique features; if $N_{k}$ is too large, then the first $N_{p}$ rows of the transformed data vector will only have the highest $k$ nodes nulled out, and the method will essentially be equivalent to discarding $\ell$-modes that have more than a fraction tol contribution from $k>k_{\rm max}$. We have found $N_{k}=\mathcal{O}(100)$ works well.
@@ 3.2 BNT
The BNT transform (Bernardeau et al., 2014) is a linear transform which reorganises the cosmic shear signal by separating the contribution from foreground structure at different redshifts. The transformed lensing kernels are narrower and as such the resulting power spectra has a closer mapping between $k$ and $\ell$. All information is preserved since the BNT transform is linear and invertible. This method benefits from a large number of tomographic bins and therefore has had limited impact of current Stage III surveys (Taylor et al., 2021) but has been shown to be able to significantly improve constraining power for a Euclid-like survey (Gu et al., 2025).
The BNT transform reorganises the cosmic shear data vector $C_{\ell,\gamma\gamma}^{ij}$ to a new data vector $\hat{C}_{\ell,\gamma\gamma}^{ab}$, where $a$ and $b$ refer to the new lensing window functions given by
$\hat{q}^{a}(\chi)=\sum_{i=1}^{N}f^{a}_{i}q^{i}(\chi)\,$, (6)
where $N$ is the number of tomographic bins and $f_{i}^{a}$ are the BNT transform coefficients. The transformation only acts on the tomographic bin space (mapping indices $i$ to transformed indices $a$), leaving the angular space unchanged.
@@
The advantage of the BNT transform is that it operates directly on the shear signal, and hence can be useful for any statistics based on the shear signal, not just the two-point statistics. It should be noted, however, that simply re-weighting the maps is not sufficient as it couples the noise, which is problematic for some high-order statistics. Methods have been developed to account for this, by applying the BNT method, smoothing the maps and then transforming back to the original basis as implemented in Euclid Collaboration et al. (2025). Another advantage is that the transformation is linear and invertible, and hence lossless.
After transformation, the mapping between angular space and 3D physical space is much cleaner in the power spectrum. Subsequent cuts in $\ell$ on the transformed power spectra can hence provide a cleaner separation of non-linear scales (Taylor et al., 2018; Deshpande et al., 2020; Taylor et al., 2021). The performance of this method is most effective when many tomographic bins are available. An example of the impact of the BNT transformation on the lensing kernels is described in Section 4.
@@ 3.3 Cross-correlation
CMB studies have used cross-correlations with foreground galaxy samples to remove the lensing contribution from the matter distribution at the redshifts of these galaxies (e.g., Qu et al. 2023). The aim in those analyses was to isolate lensing due to structure at a higher redshift, but its application to cosmic shear holds potential benefits by removing the low redshift contribution. At a fixed angular scale, it is the baryon suppression at low redshifts (equivalent to small physical scales) that contaminates the cosmic shear signal from higher redshift (equivalent to larger, more linear scales). Tracers at lower redshift can hence be used to partially remove the contribution to the shear signal from these low-redshift, small-scale structures.
@@ 4 Mock Data
The cross-correlation method relies on a low-redshift lens sample used to ‘clean’ a higher redshift source sample. When comparing the performance of this method against other methods that use only the source sample, we must be careful to account for the additional information brought by the lens density field. To perform as close an ‘apples-to-apples’ comparison as possible, we fix the total (i.e., lens + source) sample and assume that the lowest two source redshift bins can be repurposed as lens bins. The third source bin has strong overlap with neighbouring bins and is not expected to benefit significantly from the cross-correlation nulling, so we discard it. This configuration of two low-redshift lens bins and three high-redshift source bins should be roughly optimal for this form of nulling, but we caution against over-interpretation of the comparison with the other nulling methods. As the cross-correlation method has not been considered in the context of 3x2pt samples before, we consider this setup sufficient as a first test of its performance. The lensing kernels shown in Figure 2, for the standard case in the top panel and the BNT-transformed case in the lower panel.
@@
Baryon feedback is included using HMCode2020 (Mead et al., 2021) via the parameter $\Theta_{\rm AGN}$ which modulates the strength of the baryon feedback suppression. This is not a physical parameter but is related to an AGN heating parameter in the BAHAMAS simulations (McCarthy et al., 2017), defined as as $\Theta_{\rm AGN}=\log_{10}(\Delta T_{\rm heat}/{\rm K})$, which HMCode2020 was designed to be able to recover. We use $\Theta_{\rm AGN}=7.8$ which matches the fiducial BAHAMAS simulation – the ratio of the matter power spectrum with and without baryon feedback is shown in Figure 3 for a range of redshift values.
Figure 3: Suppression of the matter power spectrum due to baryon feedback ($\Theta_{\rm AGN}=7.8$) for a range of redshift values probed by the mock cosmic shear data in this analysis. Previous studies have shown that while suppression of the matter power spectrum due to baryons increases with decreasing redshift, it only evolves weakly for $z\lesssim 1$, as shown here (van Daalen et al., 2011; McCarthy et al., 2017). The evolution is non-monotonic with redshift due to different effects being more or less dominant at different times, for example gas cooling versus feedback processes (Chisari et al., 2018).